# Gravitational wave probes of particle dark matter: a review

### v1

@@ I Introduction
While the observational evidence indicates that DM may exist, it does not explain what it is made of. Could it be a single new particle, such as the axion Peccei and Quinn (1977a, b); Weinberg (1978) or dark photon Galison and Manohar (1984); Holdom (1986); Pierce et al. (2018); Filippi and De Napoli (2020), a collection of new particles, or macroscopic, in the form of PBHs formed within $\sim 1$ second of the Big Bang Green and Kavanagh (2021)? Or, particle DM may (or may not) exist alongside PBHs, which may comprise a portion or the totality of DM Hawking (1971); Carr et al. (2010); Clesse and García-Bellido (2015); Carr et al. (2016); Clesse and García-Bellido (2017); Clesse and Garcia-Bellido (2022); Miller (2024).
@@
The numerous hypotheses regarding the existence of DM suggest that the mass of its constituents can vary across hundreds of orders of magnitude Belenchia et al. (2022). Significant time and resources, however, have been spent on searching a somewhat narrow mass range comprising so-called Weakly Interacting Massive Particles (WIMPs) Jungman et al. (1996). WIMPs have been probed through collider experiments Goodman et al. (2010) and DM/ nucleus interactions by experiments such as CDMS Akerib et al. (2003, 2006); Ahmed et al. (2009), SuperCDMS Agnese et al. (2019), XENON Angle et al. (2008, 2011); Aprile et al. (2012, 2016, 2024), and LUX Akerib et al. (2013, 2016). Other experiments have tried to detect annihilating DM indirectly through its gamma-ray emission in the Galactic Center, galaxy clusters, or other areas, e.g. MAGIC Aleksic et al. (2012); Aleksić et al. (2016), VERITAS Holder et al. (2009), HESS Aharonian et al. (2006) and Fermi Hooper and Goodenough (2011), or through cosmic rays, e.g. AMS Aguilar et al. (2007, 2013); Bergstrom et al. (2013) or Fermi-LAT Abdo et al. (2009); Ackermann et al. (2012). Despite these efforts, however, WIMPS remain undetected [^1] DAMA claims to see a periodic oscillation of a DM signal, though this result has not yet been reproduced Petriello and Zurek (2008), and so it is worthwhile to ask whether other models could better explain DM. Thus, the experimental development from the particle physics community has widened to include probing ultralight DM through torsion balance experiments, such as Eöt-Wash Su et al. (1994); Schlamminger et al. (2008) and the MICROSCOPE satellite Touboul et al. (2012); Bergé et al. (2018), whose results can be interpreted as constraints on DM, and resonant cavity experiments, such as ADMX Du et al. (2018), among others. It is therefore also worth asking whether high-precision GW interferometers, such as LIGO, Virgo and KAGRA Aasi et al. (2015); Acernese et al. (2015); Aso et al. (2013), can be used to search for DM, and in particular, whether continuous gravitational-wave (CWs) can be used as the probe. While not yet as extensive as particle physics experiments, GW probes of DM have already, in many cases, surpassed limits from existing experiments, or will beat them as the detectors, both in ground and in space, develop.
@@
CWs are quasi-monochromatic, quasi-infinite-duration signals that canonically arise from asymmetrically rotating, lumpy neutron stars. Over the last few decades, much research has focused on developing ways to probe such neutron stars, both those that are known electromagnetically and those that may only be emitting GWs (“gravitars”) Sieniawska and Bejger (2019); Tenorio et al. (2021); Riles (2023); Piccinni (2022); Miller (2023); Wette (2023). Though the signal model is simple, the unknown sky positions and the uncertain neutron-star physics (e.g. the unknown equation of state, “spin wandering” Akmal et al. (1998); Cutler (2002), changes in the magnetosphere Lyne et al. (2010), etc.) complicates these searches, and thus methods that are robust not only to non-Gaussian noise disturbances but also to theoretical uncertainties had to be developed to look for such systems Krishnan et al. (2004); Dhurandhar et al. (2008); Astone et al. (2014); Suvorova et al. (2016); Bayley et al. (2019). Until now, such searches have only yielded constraints on the maximum possible deformation of the size of known Abbott et al. (2021a, 2022a); Ashok et al. (2021, 2024); Abac et al. (2025) and unknown neutron stars Abbott et al. (2022b); Steltner et al. (2023).
@@
It turns out that with CWs, we can not only probe neutron stars, but also a wide range of DM models. If DM particles couple to standard model ones, we could see them in GW data via their interactions with detector components, which would give rise to a narrow-band, stochastic “correlated-noise” signal Pierce et al. (2018); Guo et al. (2019); Grote and Stadnik (2019); Ismail et al. (2023). Separately, if a new boson that may or may not be DM exists and couples to gravity, and is of just the right mass, it will form clouds around rotating BHs, extracting mass and spin from them as they grow, and emitting CWs as the DM particles annihilate Brito et al. (2015). Heavy, strongly-interacting DM could induce collapses of stars that form in binaries, and as these objects orbit each other, they will emit GWs at particular frequencies distinct from if they were ordinary stars of the same mass Singh et al. (2021); Bhattacharya et al. (2024). Such ultralight particles could also interact with the material inside neutron stars, changing their moments of inertia and thus their CW emission Khmelnitsky and Rubakov (2014). CWs can be used to probe the millisecond pulsar hypothesis to the GeV excess, which, as it is more constrained, lends more strength to the WIMP DM annihilation hypothesis for the GeV excess Calore et al. (2019); Miller and Zhao (2023); Bartel and Profumo (2024). Finally, axions can be constrained using the very slow orbit of the Hulse-Taylor binary and other known binaries to be below to be $f_{a}<\mathcal{O}(10^{11})$ GeV, depending on the system Kim and Yang (2024).
@@ III.4 Pulsar timing arrays
Ultralight scalar DM will also cause deviations in the times of arrivals of pulses as given in Khmelnitsky and Rubakov (2014), though the mass of such particles are 10 orders of magnitude lower than those probed with ground-based interferometers.
@@ III.4.1 Gravitationally-interacting ultralight dark matter
As discussed before, DM could interact only via gravity, and alter the distances between the pulsars and earth, analgous to GWs. Using European pulsar timing array data, a search was performed in the residuals for this kind of DM interaction. In Fig. 8, we show recent constraints on ultralight scalar DM in two ways: (1) the minimum detectable strain amplitude as a function of DM mass, and (2) the fraction of DM that ultralight dark matter could compose. We can see that the masses below $\sim 10^{-23.2}$ eV can be well constrained by pulsar timing arrays, and that ultralight dark matter cannot make up all of DM at these masses. Fig. 8 presents upper limits in three scenarios: (1) uncorrelated, (2) correlated and (3) pulsar-correlated.
@@ III.4.3 Constraints on vector dark matter
Fig. 11 shows constraints on the coupling constants $g_{B}=\epsilon_{B}$ and $g_{B-L}=\epsilon_{B-L}$ as a function of the DM mass and oscillation frequency. Again, the gray-shaded region indicates the regime in which the coupling of DM to gravity becomes stronger than the coupling of DM to standard-model particles. We can see that the NANOGrav constraints greatly surpass those that come from MICROSCOPE at very low masses. Furthermore, these constraints assume that vector bosons constitute all of DM, which may not be true. If they only comprise a fraction $f$ of DM, then the constraints will be weakened by $\sqrt{f}$, because of the $\sqrt{\rho_{\text{DM}}}$ term in the amplitude of these interaction signals. Furthermore, the stochastic GW background w
@@ V GWs from ultralight bosonic clouds around rotating BHs
It has been shown, under a wide range of assumptions of BH populations, spins, redshift distributions, etc., that such systems would be detectable by CW searches in advanced LIGO, Virgo and KAGRA data and/or in future detectors, such as ET and Cosmic Explorer Arvanitaki et al. (2015b); Zhu et al. (2020). Furthermore, both galactic BHs Palomba et al. (2019) and remnants of mergers of BHs outside the galaxy could also be detected, the latter of which, in the scalar case, only being visible in ET and Cosmic Explorer Arvanitaki et al. (2010, 2015b); Isi et al. (2019). Annihilating vector boson cloud systems, however, emit stronger, though shorter, GW signals Brito et al. (2015); Baryakhtar et al. (2017); Siemonsen et al. (2023b); Jones et al. (2023, 2024), and potentially may contain electromagnetic counterparts if the vector boson kinetically mixes with the ordinary photon Siemonsen et al. (2023a).
@@
Note that while spin-0 and spin-1 boson clouds are relatively well understood theoretically, spin-2 clouds are much more complicated. The superradiant instability of spin-2 fields occurs much faster than the others East and Siemonsen (2023), and since the backreaction of the instability has not been computed due to difficulties arising from a nonlinear coupling between the spin-2 particle and gravity, it is not yet known whether the observational signatures are similar to the vector or scalar cases, and whether there is a dependency on the specific nonlinear theory chosen.
@@ V.4 Clouds in binary black hole systems
We know that BHs form binaries, and if ultralight dark matter exists in the form of ultralight bosons, it should imprint some signature in the GWs arising from inspiraling binary black holes Baumann et al. (2019a). The GW signal will be modified due to tidal disruptions induced by the companion, and the multiple moments of the cloud Baumann et al. (2019a, b, 2020). These kinds of signals could be visible in future space-based detectors, since the deviations primarily occur in the early inspiral stages of the system, i.e. at frequencies much lower than currently accessible by ground-based GW interferometers. In both cases, observations of the inspiral of binary systems could indicate the presence of boson clouds around one or both of the objects. Additionally, once the binary reaches a certain stage of its evolution, the impact of both the spin-induced quadrupole moment and tidal forces will lessen, since the clouds will become disrupted by their companions. If we can observe variations in both the spin-induced quadrupole moment, parameterized as $\kappa$, and the tidal deformability, parameterized as $\Lambda$, over the course of the lifetime of the binary, it would indicate the presence of a (disrupted) boson cloud Baumann et al. (2019a). More recently, it has been show that the boson cloud could be “ionized” as the orbital separation approaches the size of the cloud. Self-interactions could also play a role in the evolution of a binary, in which two types of clouds can actually coexist around the rotating BH through mode coupling Boudon et al. (2024); Aurrekoetxea et al. (2024a); Takahashi et al. (2024). We will now describe different ways in which boson clouds could affect GWs from binary black hole systems.
@@ V.4.1 Spin-induced quadrupole moment
Ref. Krishnendu et al. (2017) showed that $\kappa$ can be used as a distinguishing parameter for different kinds of objects, and also a test of the no-hair theorem: as an example, $\kappa=1$ for BHs Hansen (1974), while $\kappa\sim[1.4,8]$ for neutron stars depending on the equation-of-state Laarakkers and Poisson (1999); Pappas and Apostolatos (2012). In the case of boson cloud systems, $\kappa$ will take on vastly different values, of $\mathcal{O}(1000)$, depending on the boson cloud and BH masses, and the extent to which the Kerr metric is altered in the presence of a cloud with enough mass compared to the BH.
@@ VI.2.2 Viterbi
Recently, the Viterbi algorithm has also been adapted to search for GWs from annihilating vector boson clouds, which have timescales significantly shorter than in the scalar case Jones et al. (2023, 2024), and for which numerical waveforms must be used to gauge sensitivity Siemonsen and East (2020); Siemonsen et al. (2023b).

### v3

@@ I Introduction
While the observational evidence indicates that DM may exist, it does not explain what it is made of. Could it be a single new particle, such as the axion Peccei and Quinn (1977a, b); Weinberg (1978) or dark photon Galison and Manohar (1984); Holdom (1986); Pierce et al. (2018); Filippi and De Napoli (2020), a collection of new particles, or macroscopic, in the form of PBHs formed within $\sim 1$ second of the Big Bang Green and Kavanagh (2021)? Or, particle DM may (or may not) exist alongside PBHs, which may comprise a portion or the totality of DM Hawking (1971); Carr et al. (2010); Dai et al. (2009); Clesse and García-Bellido (2015); Carr et al. (2016); Clesse and García-Bellido (2017); Clesse and Garcia-Bellido (2022); Miller (2024).
@@
The numerous hypotheses regarding the existence of DM suggest that the mass of its constituents can vary across hundreds of orders of magnitude Belenchia et al. (2022). Significant time and resources, however, have been spent on searching a somewhat narrow mass range comprising so-called Weakly Interacting Massive Particles (WIMPs) Jungman et al. (1996). WIMPs have been probed through collider experiments Goodman et al. (2010) and DM/ nucleus interactions by experiments such as CDMS Akerib et al. (2003, 2006); Ahmed et al. (2009), SuperCDMS Agnese et al. (2019), XENON Angle et al. (2008, 2011); Aprile et al. (2012, 2016, 2024), and LUX Akerib et al. (2013, 2016). Other experiments have tried to detect annihilating DM indirectly through its gamma-ray emission in the Galactic Center, galaxy clusters, or other areas, e.g. MAGIC Aleksic et al. (2012); Aleksić et al. (2016), VERITAS Holder et al. (2009), HESS Aharonian et al. (2006) and Fermi Hooper and Goodenough (2011), or through cosmic rays, e.g. AMS Aguilar et al. (2007, 2013); Bergstrom et al. (2013) or Fermi-LAT Abdo et al. (2009); Ackermann et al. (2012). Despite these efforts, however, WIMPS remain undetected [^1] DAMA claims to see a periodic oscillation of a DM signal, though this result has not yet been reproduced Petriello and Zurek (2008), and so it is worthwhile to ask whether other models could better explain DM. Thus, the experimental development from the particle physics community has widened to include probing ultralight DM through (1) torsion balance experiments, such as Eöt-Wash Su et al. (1994); Schlamminger et al. (2008) and the MICROSCOPE satellite Touboul et al. (2012); Bergé et al. (2018), whose results can be interpreted as constraints on DM, (2) resonant cavity experiments, such as ADMX Du et al. (2018), (3) the Event Horizon telescope Doeleman et al. (2009); Akiyama et al. (2019), which could measure photons lensed by boson clouds and thus affect BH images Chen et al. (2023), and which could measure changes in the polarization of the electric field when the light passes through the axion background Chen et al. (2022, 2022), and (4) astrometry experiments, such as Gaia Prusti et al. (2016) and the Roman Space Telescope Akeson et al. (2019), that can meausre stochastic metric perturbations sourced by ultralight dark matter Kim (2024). It is therefore also worth asking whether high-precision GW interferometers, such as LIGO, Virgo and KAGRA Aasi et al. (2015); Acernese et al. (2015); Aso et al. (2013), can be used to search for DM, which has also been briefly reviewed in Bertone et al. (2020).
Axions and other forms of DM could be probed via their interaction with binary neutron star inspirals that could be detected in the future with LIGO, Virgo and KAGRA Hook and Huang (2018); Croon et al. (2018). The observational signature of axions on binary neutron star inspirals would be a dephasing effect induced by each object having a scalar charge, and the mass of the scalar field Hook and Huang (2018); Zhang et al. (2021). Furthermore, a variety of follow-up studies Sagunski et al. (2018); Huang et al. (2019) have been performed regarding the detectability of the effects of axions on binary neutron star inspirals, with Ref. Zhang et al. (2021) placing the first-ever constraints on the nuclear coupling of axion-like particles with masses below $10^{-11}$ eV using the detection of GWs from the binary neutron star GW170817, effectively ruling out axion decay constants of [$1.6\times 10^{16}$,$10^{18}$] GeV. Furthermore, even the very slow orbit of the Hulse-Taylor binary and other known binaries constrains the axion decay constant to be below certain values depending on the system Kumar Poddar et al. (2020) to be $f_{a}<\mathcal{O}(10^{11})$ GeV, and resonances could appear during the orbit when the orbital frequency of the binary matches that of the axion Kadota et al. (2024); Kim and Yang (2024). Finally, axions and the DM energy density can also be constrained via the evidence for stochastic GW background from pulsar timing arrays Guo et al. (2024a); Chen et al. (2024a) and also explain the mass gap in observed BHs Croon and Sakstein (2023).
@@
In particular, in this review, we would like to explore how continuous gravitational-wave (CWs) can be used as a probe of DM. CWs are quasi-monochromatic, quasi-infinite-duration signals that canonically arise from asymmetrically rotating, lumpy neutron stars. Over the last few decades, much research has focused on developing ways to probe such neutron stars, both those that are known electromagnetically and those that may only be emitting GWs (“gravitars”) Sieniawska and Bejger (2019); Tenorio et al. (2021); Riles (2023); Piccinni (2022); Miller (2023); Wette (2023). Though the signal model is simple, the unknown sky positions and the uncertain neutron-star physics (e.g. the unknown equation of state, “spin wandering” Akmal et al. (1998); Cutler (2002), changes in the magnetosphere Lyne et al. (2010), etc.) complicates these searches, and thus methods that are robust not only to non-Gaussian noise disturbances but also to theoretical uncertainties had to be developed to look for such systems Krishnan et al. (2004); Dhurandhar et al. (2008); Astone et al. (2014); Suvorova et al. (2016); Bayley et al. (2019). Until now, such searches have only yielded constraints on the maximum possible deformation of the size of known Abbott et al. (2021a, 2022a); Ashok et al. (2021, 2024); Abac et al. (2025) and unknown neutron stars Abbott et al. (2022b); Steltner et al. (2023).
@@
It turns out that with CWs, we can not only probe neutron stars, but also a wide range of DM models. If DM particles couple to standard model ones, we could see them in GW data via their interactions with detector components, which would give rise to a narrow-band, stochastic “correlated-noise” signal Pierce et al. (2018); Guo et al. (2019); Grote and Stadnik (2019); Ismail et al. (2023). Separately, if a new boson that may or may not be DM exists and couples to gravity, and is of just the right mass, it will form clouds around rotating BHs, extracting mass and spin from them as they grow, and emitting CWs as the DM particles annihilate Brito et al. (2015). Heavy, strongly-interacting DM could induce collapses of stars that form in binaries, and as these objects orbit each other, they will emit GWs at particular frequencies distinct from if they were ordinary stars of the same mass Singh et al. (2021); Bhattacharya et al. (2024). Such ultralight particles could also interact with the material inside neutron stars, changing their moments of inertia and thus their CW emission Khmelnitsky and Rubakov (2014). CWs can be used to probe the millisecond pulsar hypothesis to the GeV excess, which, as it is more constrained, lends more strength to the WIMP DM annihilation hypothesis for the GeV excess Calore et al. (2019); Miller and Zhao (2023); Bartel and Profumo (2024).
In LABEL:fig:summary, we show a landscape of the kinds of particle DM that could be probed with current and future interferometers, both by using GWs and and through direct detection of particles interacting with the instruments. There are thus so many different kinds of DM that can be searched for by using instruments that were not strictly designed to probe the existence of DM.
@@ III.4 Pulsar timing arrays
Ultralight scalar DM will also cause deviations in the times of arrivals of pulses as given in Khmelnitsky and Rubakov (2014), though the mass of such particles are 10 orders of magnitude lower than those probed with ground-based interferometers. It can also induce stochastic perturbations in the metric, which would also induce time-varying changes in the times of arrivals of pulses Kim and Mitridate (2024), but generally could be seen in interferometers with long baselines Kim (2023).
@@ III.4.1 Gravitationally-interacting ultralight dark matter
As discussed before, DM could interact only via gravity, and alter the distances between the pulsars and earth, analogously to GWs. Using European pulsar timing array data, a search was performed in the residuals for this kind of DM interaction. In Fig. 8, we show recent constraints on ultralight scalar DM in two ways: (1) the minimum detectable strain amplitude as a function of DM mass, and (2) the fraction of DM that ultralight dark matter could compose. We can see that the masses below $\sim 10^{-23.2}$ eV can be well constrained by pulsar timing arrays, and that ultralight dark matter cannot make up all of DM at these masses. Fig. 8 presents upper limits in three scenarios: (1) uncorrelated, (2) correlated and (3) pulsar-correlated.
@@ III.4.3 Constraints on vector dark matter
Fig. 11 shows constraints on the coupling constants $g_{B}=\epsilon_{B}$ and $g_{B-L}=\epsilon_{B-L}$ as a function of the DM mass and oscillation frequency. Again, the gray-shaded region indicates the regime in which the coupling of DM to gravity becomes stronger than the coupling of DM to standard-model particles. We can see that the NANOGrav constraints greatly surpass those that come from MICROSCOPE at very low masses. Furthermore, these constraints assume that vector bosons constitute all of DM, which may not be true. If they only comprise a fraction $f$ of DM, then the constraints will be weakened by $\sqrt{f}$, because of the $\sqrt{\rho_{\text{DM}}}$ term in the amplitude of these interaction signals.
@@
The recent NANOGRAV evidence for a stochastic GW background is also consistent with a vector ultralight dark matter hypothesis. Ref. Chowdhury et al. (2024) has found that ultralight dark matter that couples to muons ($L_{\mu}-L_{\tau})$ is actually more consistent with NANOGRAV’s observations than a stochastic GW background. This type of ultralight dark matter would cause the pulsars to oscillate (because they are composed of muons) and induce a spin-dependent frequency shift in the pulses that would change the angular correlations with respect to the stochastic GW background case (because of the gravitational potential of the ultralight dark matter). The mass and coupling constants of this ultralight dark matter are not ruled out by any other experiment, and lie in the region
@@ V GWs from ultralight bosonic clouds around rotating BHs
It has been shown, under a wide range of assumptions of BH populations, spins, redshift distributions, etc., that such systems would be detectable by CW searches in advanced LIGO, Virgo and KAGRA data and/or in future detectors, such as ET and Cosmic Explorer Arvanitaki et al. (2015b); Zhu et al. (2020). Furthermore, both galactic BHs Palomba et al. (2019) and remnants of mergers of BHs outside the galaxy could also be detected, the latter of which, in the scalar case, only being visible in ET and Cosmic Explorer Arvanitaki et al. (2010, 2015b); Isi et al. (2019). Annihilating vector boson cloud systems, however, emit stronger, though shorter, GW signals Brito et al. (2015); Baryakhtar et al. (2017); Siemonsen et al. (2023b); Jones et al. (2023, 2025), and potentially may contain electromagnetic counterparts if the vector boson kinetically mixes with the ordinary photon Siemonsen et al. (2023a).
@@
Note that while spin-0 and spin-1 boson clouds are relatively well understood theoretically, spin-2 clouds are much more complicated. The superradiant instability of spin-2 fields occurs much faster than the others East and Siemonsen (2023), and since the backreaction of the instability has not been computed due to difficulties arising from a nonlinear coupling between the spin-2 particle and gravity, it is not yet known whether the observational signatures are similar to the vector or scalar cases, and whether there is a dependency on the specific nonlinear theory chosen. Such ultralight bosons could also be used to probe the quantum nature of BHs Mitra et al. (2024).
@@ V.4 Clouds in binary black hole systems
We know that BHs form binaries, and if ultralight dark matter exists in the form of ultralight bosons, it should imprint some signature in the GWs arising from inspiraling binary black holes Baumann et al. (2019a). The GW signal will be modified due to tidal disruptions induced by the companion, the multiple moments of the cloud Baumann et al. (2019a, b, 2020) or through a dynamical friction drag forced induced by the ultralight dark matter. These kinds of signals could be visible in future space-based detectors, since the deviations primarily occur in the early inspiral stages of the system, i.e. at frequencies much lower than currently accessible by ground-based GW interferometers. In both cases, observations of the inspiral of binary systems, and in particular extreme mass ratio inspiral (EMRI) systems Zhang and Yang (2019); Brito and Shah (2023), could indicate the presence of boson clouds around one or both of the objects if there is some dephasing with respect to the vacuum . Additionally, once the binary reaches a certain stage of its evolution, the impact of both the spin-induced quadrupole moment and tidal forces will lessen, since the clouds will become disrupted by their companions. If we can observe variations in both the spin-induced quadrupole moment, parameterized as $\kappa$, and the tidal deformability, parameterized as $\Lambda$, over the course of the lifetime of the binary, it would indicate the presence of a (disrupted) boson cloud Baumann et al. (2019a). More recently, it has been show that the boson cloud could be “ionized” as the orbital separation approaches the size of the cloud. Self-interactions could also play a role in the evolution of a binary, in which two types of clouds can actually coexist around the rotating BH through mode coupling Boudon et al. (2024); Aurrekoetxea et al. (2024a); Takahashi et al. (2024). We will now describe different ways in which boson clouds could affect GWs from binary black hole systems.
@@ V.4.1 Spin-induced quadrupole moment
Ref. Krishnendu et al. (2017) and Ref. Rahman and Bhattacharyya (2023) showed that $\kappa$ can be used as a distinguishing parameter for different kinds of objects, and also a test of the no-hair theorem: as an example, $\kappa=1$ for BHs Hansen (1974), while $\kappa\sim[1.4,8]$ for neutron stars depending on the equation-of-state Laarakkers and Poisson (1999); Pappas and Apostolatos (2012). In the case of boson cloud systems, $\kappa$ will take on vastly different values, of $\mathcal{O}(1000)$, depending on the boson cloud and BH masses, and the extent to which the Kerr metric is altered in the presence of a cloud with enough mass compared to the BH.
@@ VI.2.2 Viterbi
Recently, the Viterbi algorithm has also been adapted to search for GWs from annihilating vector boson clouds, which have timescales significantly shorter than in the scalar case Jones et al. (2023, 2025), and for which numerical waveforms must be used to gauge sensitivity Siemonsen and East (2020); Siemonsen et al. (2023b).